OpenStudy (anonymous):

PLEASE HELP!! WILL MEDAL BEST ANSWER! Please help me! If g(x) = x2 + 3, find g(4) Help??

OpenStudy (aliqanber):

g(4)= 4^2 + 3 = 19

OpenStudy (malijhaa):

Substitute the values. 3. g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there?

OpenStudy (anonymous):

Oh ok, I see. I saw these three answer before and I am not sure which one is realy true: 19, 11?

OpenStudy (anonymous):

two answers**

OpenStudy (anonymous):

Should I go with 19???

OpenStudy (anonymous):

I GOT IT RIGHT!! The answer is certainly 19, and thank you!

alones (alones):

Nice job ;)

OpenStudy (anonymous):

Yup!